# Searching for MeV-scale Axion-like Particles and Dark Photons with PandaX-4T Tao Li

### v1

@@
Searching for MeV-scale Axion-like Particles and Dark Photons with PandaX-4T
@@ Abstract
Axion-like particles (ALPs) and dark photons (DPs) are viable dark matter particle candidates.
We have searched for possible ALP/DP signals in the PandaX-4T liquid xenon detector using 94.8 days of data.
@@
No signal excess over background expectations is observed, and we have established the most stringent exclusion limits for most ALP/DP masses ranging from 150 keV/ to 1 MeV/.
@@
Among these models, the axion-like particles (ALPs) and dark photons (DPs), also referred to as bosonic super-WIMPs [16, 17, 18], are of experimental interest.
@@
The absorption cross-section equals to [18],
in which () is the rest mass of the ALP (electron), is the fine-structure constant, is the velocity of the incoming ALP, is the dimensionless coupling constant between the electrons and ALP, and is the photoelectric effect cross-section for a photon with an energy of .
Similarly, the cross-section for DP is
@@
Among the searches of ALPs and DPs [19, 20, 21, 22, 23, 24, 25, 26, 27, 28], XENONnT [20] is leading the limit at the masses below 140 keV/, while GERDA [21] and COSINE-100 [22] have set the most significant constraints in the
@@
In this paper, we use the commissioning data set (Run0) of the PandaX-4T experiment, a total of 94.8 days of data from November 28, 2020, to April 16, 2021, to search for ALP and DP signals.
The targeted ALP or DP masses are between 30 keV/ and 1 MeV/, while the search is performed in an energy region of interest (ROI) of 25 keV to 1 MeV.
@@
The active volume containing 3.7 tonnes of natural xenon is surrounded by a field cage with an anode on the top and a cathode on the bottom.
@@
The total energy of an event is calculated by combining and collected by the bottom PMT array, according to the formula [34].
PDE, EEE, and are the photon detection efficiency for , electron extraction efficiency, and the single-electron gain for , respectively.
The horizontal position is obtained based on the maximum likelihood estimation with the observed charge distribution in the top PMT array and the photon acceptance functions derived from optical Monte Carlo (MC) simulations.
@@
This determines the fiducial mass (FM) to be kg by scaling the percentage of [30], corresponding to the total exposure of kg yr.
The detector is calibrated using external radioactive 137Cs, 60Co, and 232Th sources.
The non-uniformity of spatial energy response is monitored and corrected using the internal calibration source Kr.
Compared with the previous analysis [30], the energy reconstruction has been improved in three aspects.
The temporal variation of energy response is characterized by the signals of 222Rn progenies and corrected accordingly.
The linearity of energy reconstruction is further improved by adding one more high-statistics 41.5 keV peak from Kr, in additional to Xe, Xe, 127Xe, 60Co, 40K, and 232Th.
Consequently, we have improved the energy resolution, specifically from 3.6% to 3.0% at 208 keV, which allows improvements in estimating activities of short-lived xenon isotopes mentioned later.
@@
represents the 5-parameter detector response model (see text), with means and uncertainties determined from calibration fit.
The energy response is modeled with five parameters.
The energy resolution is modeled as a Gaussian function with the width constructed as , with energy in the unit of keV.
The energy scale is defined as to account for possible bias with respect to the reconstructed energy .
The measured energy spectrum is a convolution of the true energy spectrum with the five-parameter response model.
The parameters and uncertainties are determined by fitting the peaks of 41.5 keV (Kr), 164 keV (Xe), 236 keV (Xe and Xe) from the calibration data, and 1460 keV (40K) outside the ROI from Run0, therefore completely uncorrelated with the final search fit.
The extracted values and uncertainties of the parameters (Table 1) will be used as priors, together with the covariance matrix in fitting the Run0 data.
The total detection efficiencies for signals of ALPs/DPs are the product of data quality cut efficiency, SS cut efficiency, and ROI acceptance.
The adjusted cut criteria are validated using calibration data and subsequently applied to the entire Run0, resulting in an efficiency of ()%.
@@
The ratio of the number of SS and SS+MS events within the ROI is calculated using BambooMC, a Geant4-based Monte Carlo (MC) framework [35], and verified through 232Th calibration data.
The SS/(SS+MS) ratio of 232Th calibration data in the ROI is %, with the uncertainty from the difference between data and simulation averaged over the energy range.
The relative systematic uncertainty from 232Th calibration data is conservatively used for both the signal and background for the SS fraction.
On top of the SS fraction, the quality cut efficiency is energy dependent, ranging from 100.0% to 95.7% in the ROI based on MC.
The ROI acceptance for signals is close to 100%.
The liquid xenon density and FM are two other systematic factors for ALP/DP detections.
The liquid xenon density is inherited from Ref. [29] with a relative uncertainty of 0.13%.
The relative difference between the FM calculated from geometry and that derived from Kr events distribution is 1.57%, which is taken as the systematic uncertainty.
@@
The Kr/Xe concentration is parts per trillion, assuming an isotopic abundance of for Kr [36].
The 212Pb activity of is determined from the rate of 212Po [30].
The 214Pb rate is left float in the fit, the same as the previous analyses [29, 30].
The energy spectrum of the elastic scattering of solar and 7Be neutrinos on electrons is adopted from the Ref. [37], with an uncertainty of approximately 10% [38], resulting in events within ROI.
Figure 1: Time evolution of 127Xe, Xe and Xe in SS spectrum during the Run0 of PandaX-4T experiment.
The starting time of the axis is defined as the beginning of the dataset.
@@
Table 2: The background contribution in the ROI. The fitted counts are from the background only fit.
@@
Figure 2: Fit of background and a hypothetical 990 keV/ DM signal to SS data from 25 keV to 1050 keV with bin size of 1 keV.
The x-axis represents the reconstructed energy in the data.
@@
The contributions from 136Xe and 124Xe double electron capture are calculated based on the half-life measurements in PandaX-4T [29, 39].
The expected events in ROI are and , respectively.
@@
Xe and Xe results in mono-energetic peaks at 164 keV and 236 keV, with the half-life of 11.8 days and 8.9 days, respectively.
125Xe, with a half-life of 16.9 hours, undergoes electron capture and decays to the relatively long-lived isotope 125I, with a half-life of 59.4 days.
Given the low statistics, 125Xe is treated as a floating parameter in the fit.
The activity of 125I is calculated based on its temporal modeling [39], resulting in events in the ROI.
133Xe is also left float in the fit since the spectrum is continuous, characterized by a 346 keV endpoint decay with corresponding de-excitation 81 keV ray.
In addition, 127Xe was introduced by a batch of approximately 30 kg of xenon from above-ground (exposed to cosmogenic neutrons), injected into the detector during the data taking.
127Xe decays via electron capture with a half-life of 36.4 days.
The de-excitation of the daughter 127I generates mono-energetic peaks at 208 keV, 236 keV, 380 keV, and 408 keV.
The estimated numbers of 127Xe, Xe, and Xe events are obtained by fitting the time evolution in the SS data, with decay half-lives also set free.
In order to increase statistics, a larger FM of 2.43 tonnes is selected.
Each peak is characterized by a Gaussian function plus a linear background.
@@
Based on the temporal fit, the estimated number of events for the 164 keV, 208 keV, 236 keV, 380 keV, and 408 keV peaks in ROI are shown in Table 2.
We employ a binned likelihood method, with the likelihood function constructed as
where and are the expected and observed numbers of events in the energy bin, respectively.
The Gaussian penalty term of the energy response contains the five-parameter (Table 1) and the covariant matrix .
The Gaussian penalty terms constrain the nuisance parameters , , and , which are the relative uncertainties of the overall efficiency, the signal selection (Table 1), and the background model (Table 2), respectively.
is the number of Gaussian-constrained nuisance paramters.
is defined as
@@
A background-only fit is performed before the signal fits, with /NDF=1.14.
The fitted background contributions are shown in Table 2 and agree with the expected values.
@@
Fitted nuisance parameters are within 1 of the input values, except for and in the energy response model, which are pulled by 1.9 and 2.5 , respectively.
This suggests that the energy resolution function derived solely by mono-energetic peaks is insufficient to describe the full measured spectrum.
We conduct a scanned fit to the SS spectrum, including Gaussian peaks of the hypothetical DM signals, with DM masses spanning from 30 keV/ to 1 MeV/ with a step size of 10 keV/.
An illustrative example of the fit results for DM mass of 990 keV/ with data-fit residuals is presented in Fig. 2.
At this mass, the majority of residuals is less than 3.
The fitted signal count is .
The local significances of two DM masses, 250 keV and 260 keV, are found to be between 3 and 4, occurring near the Gaussian background of 236 keV (Xe and Xe).
Taking into account the look elsewhere effect [41, 42], the global significance is less than 1. Therefore, no significant evidence for a signal is observed within the mass range of [30 keV/, 1 MeV/].
@@
The results from this work are represented by the red curve starting at 30 keV, while the curve below 25 keV is taken from Ref. [43].
The upper limits at 90% confidence level (C.L.) on the event rate have been set and converted to upper limits of coupling strength (Fig. 3), as described in Eq. 1.
@@
Our limits are the most competitive over a wide region of the masses, ranging from 150 keV/ to 1 MeV/, with an average improvement of 1.5 times better than the existing results [22, 21].
The coupling of ALPs (DPs) to electrons of () at the mass of 990 keV/ derived from the fit shown in Fig. 2 is 2.5 times better than the COSINE-100 result [22].
@@
It is noteworthy that the weakening of our upper limits with increasing mass is primarily due to the steep decrease of the photoelectric cross-section, as we solely searched for absorption events.
In the MeV mass regions of the ALPs and DPs, the cross-section of Compton-like process [22], which are mostly MS events, becomes more prominent.
@@
In summary, we have searched for ALPs and DPs with masses up to 1 MeV/ using 162.3 kg yr exposure of PandaX-4T Run0.
@@
Between 2021-2022, PandaX-4T completed the first science run (Run 1) with 163.5 days data taking, and has recently completed an upgrade and resumed physics data taking.
@@
This project is supported in part by grants from National Science Foundation of China (Nos. 12090060, 12090061, 12305121, U23B2070), a grant from the Ministry of Science and Technology of China (Nos. 2023YFA1606200, 2023YFA1606201, 2023YFA1606202), China Postdoctoral Science Foundation (No. 2023M744093), and Office of Science and Technology, Shanghai Municipal Government (grant No. 22JC1410100, 21TQ1400218).
We thank for the support by the Fundamental Research Funds for the Central Universities. We also thank the sponsorship from the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Hongwen Foundation in Hong Kong, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.

### v2

# Search for MeV-scale Axionlike Particles and Dark Photons with PandaX-4T Tao Li

@@
Search for MeV-scale Axionlike Particles and Dark Photons with PandaX-4T
@@ Abstract
Axionlike particles (ALPs) and dark photons (DPs) are viable dark matter particle candidates.
We have searched for possible ALP/DP signals in the PandaX-4T liquid xenon detector using 440 kgyr of data.
@@
No signal excess over background expectations is observed, and we have established the most stringent exclusion limits for most ALP/DP masses across the range of 150 keV/ to 1 MeV/.
The improvement is particularly significant within the mass range of 150-400 keV/,
with the average factor of 3.5 compared to previous results.
@@
Among these models, the axionlike particles (ALPs) and dark photons (DPs), also referred to as bosonic super-WIMPs [16, 17, 18], are of experimental interest.
@@
The absorption cross section equals to [18],
in which () is the mass of the ALP (electron), is the fine-structure constant, is the velocity of the incoming ALP, is the dimensionless coupling constant between the electrons and ALP, and is the photoelectric effect cross section for a photon with an energy of .
Similarly, the cross section for DP is
@@
Among the searches of ALPs and DPs [19, 20, 21, 22, 23, 24, 25, 26, 27, 28], XENONnT [20] has the leading limit at the masses below 140 keV/, while GERDA [21] and COSINE-100 [22] have set the most significant constraints in the
@@
Note that a strong cosmological constraint on axionlike particles (ALPs) with masses above 100 keV/ can be derived from the lifetime requirement, assuming that heavy ALPs couple exclusively to electrons or have generic couplings to other fermions [29]. However, if these assumptions about the couplings are relaxed, the resulting constraints carry significant uncertainties.
Consequently, the direct search for ALPs in this mass range remains a topic of high scientific interest [29].
In this Letter, we use the commissioning dataset (Run0) and the first scientific dataset (Run1) of the PandaX-4T experiment, covering the periods from November 28, 2020, to April 16, 2021, and November 16, 2021, to May 15, 2022, respectively, to search for ALP and DP signals. The durations of Run0 and Run1 are 94.8 day and 163.5 day, respectively.
The targeted ALP or DP masses are between 30 keV/ and 1 MeV/, while the search is performed in an energy region of interest (ROI) of 25 to 1050 keV.
@@
The active volume containing 3.7 ton of natural xenon is surrounded by a field cage with an anode on the top and a cathode on the bottom.
@@
To avoid systematic effects due to saturation of the top PMTs, the total energy of an event is calculated by combining the charges of and , collected by the bottom PMT array, according to the formula [35].
The detector parameters (, ) are prefitted by the mono-energy electronic recoil peaks [36].
The horizontal position is obtained based on the maximum likelihood estimation with the desaturated charge pattern of in the top PMT array and the photon acceptance functions derived from optical Monte Carlo (MC) simulations.
@@
The same fiducial volume cuts of the previous study [31] have been adopted in this analysis.
The fiducial mass (FM) is determined to be kg for Run0 and kg for Run1.
These values are obtained by scaling the percentage of , corresponding to a total exposure of 440 kgyr.
The uncertainty is determined by the LXe density, and by the difference between the geometrically calculated and rate-scaled volumes [31].
@@
denotes the five-parameter detector response model (see text), with its means and uncertainties determined from mono-energetic peaks obtained during calibration runs and outside the ROI.
The detector is calibrated using multiple mono-energetic peaks observed in the data, obtained either during dedicated calibration periods or physics runs.
The nonuniformity in the spatial energy response is monitored and corrected using the uniformly distributed 41.5 keV peak from Kr injection data.
Compared to the previous analysis [31], the energy reconstruction has been improved in two key aspects.
First, the temporal variations in light yield and charge yield are characterized using the signals from 222Rn progenies and corrected accordingly.
This correction has led to an improvement in energy resolution, e.g., from 3.6% to 3.0% at 208 keV in Run0, enabling more accurate estimation of the activities of short-lived xenon isotopes (see later).
Second, the energy response model has been further refined by incorporating the Kr peak, along with the 164 keV peak (from Xe), the 236 keV peaks (from Xe and Xe) obtained during calibration runs, and the 1460 keV peak (from 40K) outside the ROI.
These calibration data points are completely uncorrelated with those used in the final spectral fit.
The inclusion of the Kr peak allows the analysis window to be extended down to 25 keV.
The energy response is modeled using five parameters, independently for Run0 and Run1.
The energy resolution is modeled as a Gaussian function with the width constructed as , where is the reconstructed energy in the unit of keV.
The residual energy nonlinearity is modeled as where is the true energy.
The calibrated values and uncertainties (Table 1) and the covariance matrix will be used in fitting the Run0 and Run1 data.
The total detection efficiencies for signals of ALPs and DPs are the product of SS cut efficiency, data quality cut efficiency, and ROI acceptance.
@@
The ratio of SS to SS+MS events within the ROI is calculated using BambooMC, a GEANT4-based Monte Carlo (MC) framework [37], with the LXe response to ER modeled using NEST 2.0 [38], and validated through 232Th calibration data.
The SS fraction for the signal varies ranging from 100.0% to 95.7% within the ROI.
The relative systematic uncertainty from 232Th calibration data is conservatively applied on both the signal and background for the SS fraction, which is obtained from the difference between data and simulation, averaged over the energy range.
The adjusted cut criteria are validated using calibration data and subsequently applied to the entire Run0 and Run1, resulting in an efficiency of ()% and ()%, respectively.
The ROI acceptance for signal is close to 100%.
@@
The Kr/Xe concentration in Run0 (Run1) is () parts per trillion [39], assuming an isotopic abundance of for Kr [40].
The 214Pb rate is left float in the fit, the same as the previous analyses [31]. The 212Pb originated from 220Rn emanation is a subdominant background component. Because of the strong dependence of 212Pb to the circulation conditions, etc., the rate of 212Pb of each run in the fit is also set free.
The energy spectrum of the elastic scattering of solar and 7Be neutrinos on electrons is adopted from the Ref. [41], with an uncertainty of approximately 10% [42], resulting in () events within ROI in Run0 (Run1).
Figure 1: Time evolution of 127Xe, Xe and Xe in the SS spectrum during the Run0 of PandaX-4T experiment.
The starting time of the horizontal axis is defined as the beginning of the dataset.
@@
Table 2:
The total background contributions in the ROI are presented for both Run0 and Run1.
The fitted counts are obtained from the background-only fit.
@@
Figure 2:
The background-only SS data and fit, along with hypothetical 350 keV/ DM signals (ALP with , DP with ), are shown for Run0 (left) and Run1 (right) from 25 to 1050 keV, with a bin size of 1 keV.
The horizontal axis represents the reconstructed energy in the data.
@@
The contributions from 136Xe two-neutrino double beta decay () and 124Xe double electron capture are calculated based on the half-life measurements reported in PandaX-4T [30, 36].
The expected events for 136Xe within the ROI are for Run0 and for Run1, respectively.
Similarly, the expected events for 124Xe double electron capture within the ROI are for Run0 and for Run1.
@@
The neutron sources, including Am-Be and Pu-C, are used before, during, and after Run0, while no neutron calibration runs are performed during Run1.
Xe and Xe results in mono-energetic peaks at 236 and 164 keV, with the half-life of 8.9 and 11.8 day, respectively.
125Xe, with a half-life of 16.9 h, undergoes electron capture and decays to the relatively long-lived isotope 125I, with a half-life of 59.4 day.
The spectrum of 133Xe is continuous, featuring a spectrum with a 346 keV endpoint in combination with a 81 keV deexcitation ray.
In addition, 127Xe was introduced by a batch of approximately 30 kg of xenon from above-ground (exposed to cosmogenic neutrons), injected into the detector during the Run0 data taking.
127Xe decays via electron capture with a half-life of 36.4 day.
The deexcitation of the daughter 127I generates mono-energetic peaks at 208, 236, 380, and 408 keV.
The background contributions from these neutron-activated components are summarized Table 2.
Because of their clear time dependences, the expected activities of 127Xe, Xe, and Xe in Run0 are prefitted according to the time evolution in the SS data, with decay half-lives set free.
In order to increase statistics, a larger FM of 2.43 ton is selected.
To get the rate evolution, in every time division, each peak in the energy spectrum is locally characterized by a Gaussian function plus a linear background.
@@
On the other hand, the activities of 125Xe and 133Xe are determined later by the spectral likelihood fit.
For Run1, the level of 125Xe, 133Xe, and 127Xe have decayed to zero.
Small amounts of Xe and Xe in Run1 do not allow a meaningful prefit either, therefore are left float in later spectral fit.
For the activity of 125I in both runs, we adopt the results from a two-component exponential model fit in Ref. [36].
The final SS energy spectral fit is then performed using a one-dimensional binned likelihood function constructed as
where and are the expected and observed events numbers of the energy bin in the Run-, respectively.
in Run- is defined as
@@
The Gaussian penalty term of the energy response contains the five-parameter (Table 1) and the covariant matrix in Run-.
The Gaussian penalty terms are used to constrain the nuisance parameters , , and , which represent the relative uncertainties in the overall efficiency (Table 1), the signal selection (Table 1), and background model (Table 2), respectively.
The activities of 232Th, 238U, 60Co, 40K, 124Xe, 136Xe, as well as the solar and 7Be fluxes are identical in Run0 and Run1.
Other background components are treated independently for Run0 and Run1.
A background-only fit is performed prior to the signal fits, yielding a /NDF of 1.06, as illustrated in Fig. 2.
The data are consistent with the background-only model, with a -value of 0.51.
The contributions of other fitted background components, summarized in Table 2, are consistent with their expected values, except Kr which is pulled slightly upward by 1.5 .
@@
The detector response nuisance parameters () for the two runs are within 1 of their input values, except for parameter and in the energy response model of Run0 and in that of Run1, which are pulled by 2.9 , 1.8 , and 1.2 , respectively.
This suggests that the energy resolution function derived solely from mono-energetic peaks is insufficient to describe the full measured spectrum.
However, we have verified that these deviations have a negligible impact on the limits of the couplings of ALPs and DPs.
We conduct a scanned fit to the SS spectrum, including Gaussian peaks of the hypothetical DM signals.
The DM masses range from 30 keV/ to 1 MeV/ with a step size of 10 keV/.
The local significances of five DM masses are found to range between 2 and 3 .
For instance, the significance at 230 keV, which occurrs near the Gaussian background of 236 keV (from Xe and Xe), is 3.0 .
Taking into account the look elsewhere effect [44, 45], the global significance only reaches 1.5 .
Therefore, no significant evidence for a signal is observed within the mass range of [30 keV/, 1 MeV/].
@@
The results from this work are represented by the red line starting at 30 keV, while the line below 25 keV is taken from Ref. [46].
The dashed blue lines indicates the stability requirement for ALPs and DPs over the age of the Universe[29, 47].
The dotted brown line and the dot-dashed green line represent constraints derived from energy losses in red giant (RG) and horizontal branch (HB) stars, respectively [48].
The upper limits at 90% confidence level (C.L.) on the event rate have been set and converted to upper limits of coupling strength (Fig. 3), as described in Eq. (1).
@@
Our limits are the most competitive over a broad mass range, spanning from 150 keV/ to 1 MeV/, with an average improvement with a factor of 2.0 compared to existing results [22, 21].
Notably, the improvement is particularly significant in the 150-400 keV/ range, where the enhancement factor reaches 3.5.
The largest improvement (a factor of 13.4) occurs around 150 keV/, constraining the coupling of ALPs (DPs) to electrons to ().
@@
It is noteworthy that the weakening of our upper limits with increasing mass is primarily due to the steep decrease of the photoelectric cross section, as we solely searched for absorption events.
In the MeV mass regions of the ALPs and DPs, the cross section of Compton-like process [22], which are mostly MS events, becomes more prominent.
@@
In summary, we have searched for ALPs and DPs with masses up to 1 MeV/ using 440 kgyr exposure of PandaX-4T Run0 and Run1.
@@
Between 2023 and 2024, PandaX-4T underwent upgrades to its PMTs, electronics, DAQ, and external water veto systems.
The second science run (Run2) is now underway, accompanied by continuous efforts to further reduce background levels.
@@
This project is supported in part by grants from the Ministry of Science and Technology of China (No. 2023YFA1606200, No. 2023YFA1606201, No. 2023YFA1606202), National Science Foundation of China (No. 12090060, No. 12090061, No. 12305121, No. U23B2070), China Postdoctoral Science Foundation (No. 2023M744093), and Office of Science and Technology, Shanghai Municipal Government (Grants No. 22JC1410100, No. 21TQ1400218).
We are grateful for the support by the Fundamental Research Funds for the Central Universities. We also are thankful for the sponsorship from Hongwen Foundation in Hong Kong, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund.
Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.
Szydagis et al. [2011b]M. Szydagis, N. Barry,
K. Kazkaz, J. Mock, D. Stolp, M. Sweany, M. Tripathi, S. Uvarov, N. Walsh, and M. Woods, Journal of Instrumentation 6, P10002 (2011b).